# Mathematics with large language models as provers and verifiers

### v1

@@ Abstract
During 2024 and 2025 the discussion about the theorem-proving capabilities of large language models started reporting interesting success stories, mostly to do with difficult exercises (such as problems from the International Mathematical Olympiad), but also with conjectures [8] formulated for the purpose of verifying whether the artificial intelligence could prove it. In this paper we report a theorem proving feat achieved by ChatGPT by using a protocol involving different prover and verifier instances of the gpt-5 model working collaboratively. To make sure that the produced proofs do not suffer from hallucinations, the final proof is formally verified by the lean proof assistant, and the conformance of premises and conclusion of the lean code is verified by a human. Our methodology was able to solve five out of six 2025 IMO problems, and close a third of the sixty-six number theory conjectures in [4].
@@ 1 Introduction
This paper describes a protocol, involving the OpenAI Application Programming Interface (API) with a minimal human interaction, which was able to solve five out of six problems of the International Mathematical Olympiads (IMO), close twenty-two out of sixty-six open conjectures on cyclic numbers proposed in [4], as well as discover and prove several new theorems in a variety of fields of mathematics. The human intervention only arises to verify that a formal version of the generated theorem statement is consistent with its semi-formal natural language description. More precisely, we only ask the human to verify the conformance of all premises and the conclusion of the formal and the semi-formal versions of the proof [^1] At the time of publishing this technical report we are still carrying out the human part of the protocol: the proofs should not be trusted yet; but those we reported in this paper appear correct to us..
@@ 4.4 Human interaction issues
In practice, the human interaction remains very time-consuming. This is mostly due to the limited ability shown by gpt-5 in converting its own natural language proofs into formal lean codes. So, although our protocol does minimize human interaction, this interaction is remains excessive, because, aside from a handful of cases, the human has to construct the whole lean code almost from scratch. A more intensive LLM training using lean codes could improve the situation. This is another situation (as that described in Sect. 4.2) which would be better tackled by OpenAI.
@@ 5.1.2 Qualitative comparison of our system compared to others
Table 2: Number of iterations required by the TTVR loop to arrive at a solution of the IMO25 problems.
According to Table 2, the easiest problem for humans (Problem 1) is the hardest for our protocol; the second-hardest is Problem 2. We note that both problems are geometric. The benchmark found in [26] agrees on Problem 6 (no LLM-based approach could solve it), and approximately also on the relative ease with which such approaches can solve Problems 3, 4, and 5. This benchmark tells a rather different story for Problems 1 and 2: Problem 2 is much more difficult than Problem 1 — in fact it is the most difficult problem in the set 1–5. Our protocol, however, finds that the opposite holds: Problem 1 is harder than problem 2 (admittedly according to a different measure than that of the benchmark). In general, given that LLMs are based on symbolic textual training more than on geometric drawings, one might guess that LLMs find geometric problems harder to solve than other types of problems. Our tests provide one more piece of evidence to reinforce confidence in this guess.
@@ 5.2 Conjectures of primes and cyclics
Cyclic numbers are analogous to prime numbers in a certain sense, and therefore it makes sense to conjecture that many results known for primes should also apply to cyclics. The definition of a cyclic number is as follows. Let $n$ be a positive integer, and consider the cardinality $m$ of the set $C$ of positive integers $r<n$ and relatively prime to $n$ (obviously, $m$ is a function of $n$). If $n,m$ are relatively prime, then $n$ is cyclic. It is analogous to the definition of “prime” insofar as they both involve the comparison of an integer with the cardinality $m$ of a set of numbers $r<n$ that have a certain relationship with $n$ concerning integer division. If the set is that of the divisors of $n$ and $m=2$, then $n$ is prime; if the set if that of the coprimes of $n$ and $\mathrm{gcd}(n,m)=1$, then $n$ is cyclic. To render notation more compact, we introduce Euler’s totient function $\phi(n)=m$ to be the cardinality of the integers $r<n$ and coprimes with $n$, and define $n$ to be cyclic if $\mathrm{gcd}(n,\phi(n))=1$. Note that, if $n$ is prime, then $r$ is coprime with $n$ for each $r<n$, so m=n1m=n-1, and since $\mathrm{gcd}(n,n-1)=1$ for each integer $n$, each prime is also cyclic. The paper [4] goes on to list the 66 (numbered) conjectures about cyclic numbers.
Seven out of the sixty-six conjectures, namely 2, 3, 4, 36, 37, 52, 53 had been closed by C. Pomerance prior to the publication of [4] by personal communication to its author. Our system closed twenty-two of these conjectures, namely 3, 6, 9, 14, 17, 20, 32, 35, 36, 37, 41, 42, 47, 50, 51, 52, 53, 55, 56, 59, 60, 61 with seventeen proofs and five refutations (35, 50, 51, 53, 59). Among the twenty-two solved by our system, six (3, 36, 37, 52, 53, 59) had already been settled by C. Pomerance, leaving our system with 16 (previously open) conjectures closed.
conjecture 3 6 9 14 17 20 32 35 36 37 41 42 47 50 51 52 53 55 56 59 60 61
@@
Table 3: Number of iterations required by our TTVR loop to settle (Proved/Refute) the conjectures in [4].
@@ 6 Conclusion
We described a system that employs the OpenAI LLM gpt-5 in order to prove mathematical theorems. The system is based on protocol including an alternating loop of prover and verifier agents impersonated by gpt-5, followed by the creation of a formal lean code version of the proof (if any was found). A human verifies the conformance of the natural language version with the formal version of the premises and conclusion of the theorem, which ensures proof correctness. We showcase our system’s capabilities: (i) on the International Mathematical Olympiad problems in 2025, where our system was as successful as OpenAI and Google DeepMind’s internal LLMs, (ii) on the solution of 22 out of the 66 conjectures on cyclics proposed in [4], and (iii) on the solution of a list on conjectures, formulated by the system itself, in three fields of mathematics: gradient descent in machine learning, the 1-2-3 conjecture in graph theory, and sphere packings in Euclidean spaces.

### v2

@@ Abstract
During 2024 and 2025 the discussion about the theorem-proving capabilities of large language models started reporting interesting success stories, mostly to do with difficult exercises (such as problems from the International Mathematical Olympiad), but also with conjectures [8] formulated for the purpose of verifying whether the artificial intelligence could prove it. In this paper we report a theorem proving feat achieved by ChatGPT by using a protocol involving different prover and verifier instances of the gpt-5 model working collaboratively. To make sure that the produced proofs do not suffer from hallucinations, the final proof is formally verified by the lean proof assistant, and the conformance of premises and conclusion of the lean code is verified by a human. Our methodology is by no means complete or exact. It was nonetheless able to solve five out of six 2025 IMO problems, and close about a third of the sixty-six number theory conjectures in [4].
@@ 1 Introduction
This paper describes a protocol, involving the OpenAI Application Programming Interface (API) with a minimal human interaction, which was able to solve five out of six problems of the International Mathematical Olympiads (IMO), close twenty-two out of sixty-six open conjectures on cyclic numbers proposed in [4], as well as discover and prove several new theorems in a variety of fields of mathematics. The human intervention only arises to verify that a formal version of the generated theorem statement is consistent with its semi-formal natural language description. More precisely, we only ask the human to verify the conformance of all premises and the conclusion of the formal and the semi-formal versions of the proof [^1] At the time of publishing this technical report we are still carrying out the human part of the protocol: many proofs should not be trusted yet; but those we reported in this paper appear correct to us (as human verifiers)..
@@ 4.4 Human interaction issues
In practice, the human interaction remains very time-consuming. This is mostly due to the limited ability shown by gpt-5 in converting its own natural language proofs into formal lean codes. So, although our protocol does minimize human interaction, this interaction is remains excessive, because, aside from a handful of cases, the human has to construct the whole lean code almost from scratch. A more intensive LLM training using lean codes could improve the situation. This is another situation (as that described in Sect. 4.2) which would be better tackled by OpenAI. For the time being, we instructed OpenAI’s Codex interface to reduce to “axiom” any proof step that Codex can find by means of a web search. This reduces the extent of lean checking but makes it actually possible to partly fill the certification part of our protocol.
@@ 5.1.2 Qualitative comparison of our system compared to others
correct? Y Y Y Y Y NA
certified? Y NA
Table 2: Number of iterations required by the TTVR loop to arrive at a solution of the IMO25 problems, together with a human evaluation of correctness and a lean certification.
According to Table [^5] Missing entries correspond to ongoing tests. 2, the easiest problem for humans (Problem 1) is the hardest for our protocol; the second-hardest is Problem 2. We note that both problems are geometric. The benchmark found in [26] agrees on Problem 6 (no LLM-based approach could solve it), and approximately also on the relative ease with which such approaches can solve Problems 3, 4, and 5. This benchmark tells a rather different story for Problems 1 and 2: Problem 2 is much more difficult than Problem 1 — in fact it is the most difficult problem in the set 1–5. Our protocol, however, finds that the opposite holds: Problem 1 is harder than problem 2 (admittedly according to a different measure than that of the benchmark). In general, given that LLMs are based on symbolic textual training more than on geometric drawings, one might guess that LLMs find geometric problems harder to solve than other types of problems. Our tests provide one more piece of evidence to reinforce confidence in this guess.
@@ 5.2 Conjectures of primes and cyclics
Cyclic numbers are analogous to prime numbers in a certain sense, and therefore it makes sense to conjecture that many results known for primes should also apply to cyclics. The definition of a cyclic number is as follows. Let $n$ be a positive integer, and consider the cardinality $m$ of the set $C$ of positive integers $r<n$ and relatively prime to $n$ (obviously, $m$ is a function of $n$). If $n,m$ are relatively prime, then $n$ is cyclic. It is analogous to the definition of “prime” insofar as they both involve the comparison of an integer with the cardinality $m$ of a set of numbers $r<n$ that have a certain relationship with $n$ concerning integer division. If the set is that of the divisors of $n$ and $m=1$ (only $r=1<n$ divides $n$), then $n$ is prime; if the set is that of the coprimes of $n$ and $\mathrm{gcd}(n,m)=1$, then $n$ is cyclic. To render notation more compact, we introduce Euler’s totient function $\phi(n)=m$ to be the cardinality of the integers $r<n$ and coprimes with $n$, and define $n$ to be cyclic if $\mathrm{gcd}(n,\phi(n))=1$. Note that, if $n$ is prime, then $r$ is coprime with $n$ for each $r<n$, so m=n1m=n-1, and since $\mathrm{gcd}(n,n-1)=1$ for each integer $n$, each prime is also cyclic. The paper [4] goes on to list the 66 (numbered) conjectures about cyclic numbers.
Our system produced a proof or refutation for twenty-two out of the sixty-six conjectures: namely 3, 6, 9, 14, 17, 20, 32, 35, 36, 37, 41, 42, 47, 50, 51, 52, 53, 54, 56, 59, 60, 61, with seventeen proofs and five refutations (35, 50, 51, 53, 59). Seven out of the sixty-six conjectures, namely 2, 3, 4, 36, 37, 52, 53, had already been closed by C. Pomerance prior to the publication of [4], by personal communication to its author (as far as we know, Pomerance’s proofs remain unpublished to date). The five conjectures solved by Pomerance that appear in our list (3, 36, 37, 52, 53) must therefore be removed from the list of “open conjectures”. This leaves our system with 17 (previously open) conjectures for which it obtained a proof.
conjecture 3 6 9 14 17 20 32 35 36 37 41 42 47 50 51 52 53 54 56 59 60 61
@@
correct? ? Y Y N Y Y Y N ? ? Y Y Y ? ? Y Y Y Y Y
certified? ? Y Y Y Y ? ? Y ? ?
Table 3: Number of iterations required by our TTVR loop to produce either correct and lean-certified or incorrect Proofs or Refutations of the conjectures in [4].
In Table [^6] Missing entries correspond to ongoing tests. 3 we report, for each proof produced by our system: the number of TTVR iterations ($|\mbox{itn}|$), whether the conjecture was open or already closed (O/C), whether the conjecture was settled by proof or refutation (P/R), whether the authors of this paper (who are not specialists in analytic number theory) think that the semi-formal natural language proofs are correct or not, and whether the lean certificate is correct and conformant with the premises and conclusions of the corresponding proofs.
We note that some of the confirmed conjectures have been “closed” only in the asymptotic sense, not in the absolute one. For example, the proof of Conjecture 6 shows that there is an $n_{0}$ such that for every $n>n_{0}$ the conjecture holds, but we have no idea about this $n_{0}$, so Conjecture 6 is still open “for all $n\in\mathbb{N}$”. This is also the case for Conj. 32.
In one of the refuted conjectures, namely Conj. 59, our system has provided a formally correct refutation that would be unsatisfactory to any human mathematician: it addresses one of finitely many exceptions, providing no refutation for the asymptotic part.
@@ 6 Conclusion
We described a system that employs the OpenAI LLM gpt-5 in order to prove mathematical theorems. The system is based on protocol including an alternating loop of prover and verifier agents impersonated by gpt-5, followed by the creation of a formal lean code version of the proof (if any was found). A human verifies the conformance of the natural language version with the formal version of the premises and conclusion of the theorem, which ensures proof correctness. We showcase our system’s capabilities: (i) on the International Mathematical Olympiad problems in 2025, where our system was as successful as OpenAI and Google DeepMind’s internal LLMs, (ii) on the solution of about a third out of the 66 conjectures on cyclics proposed in [4], and (iii) on the solution of a list on conjectures, formulated by the system itself, in three fields of mathematics: gradient descent in machine learning, the 1-2-3 conjecture in graph theory, and sphere packings in Euclidean spaces.